OpenStudy (anonymous):

When rolling two number cubes or a set of six-sided dice, what is the probability of rolling an 11? A. 1/36 B. 1/18 C. 1/6 D. 1/3

OpenStudy (raden):

there are 2 evens of 11: (5,6) and (6,5) and all even = 6^2 = 36 so, it is 2 of 36. 2/36 = 1/18

OpenStudy (anonymous):

ohhhh wooww thanks a lot i was having no idea how to do this